OpenStudy (anonymous):

if the slope if it exist x+7y=14

OpenStudy (agreene):

if we solve for y and put in slope intercept form, we can see the slope more clearly: x+7y=14 subtract x from both sides 7y=14-x divide by 7 \[y=-\frac{14-x}{7}\] we can split this to two fractions \[y=-\frac{14}{7}+\frac{x}{7}\] simplifying and rearranging to y=mx+b form: \[y=\frac{1}{7}x-2\] so our slope is 1/7

OpenStudy (anonymous):

Slope would be -1/7 not 1/7, in the second step you wrote, \[y=-(14-x)/7\] instead of \[y=(14-x)/7\]